Suppose that every firm in an industry has identical cost curves given by

C(q) = 3q2 + 2700

MC(q) = 6q

What is the value of the efficient scale, qE ? Remember, the efficient scale occurs where MC(q) = ATC(q)
Answer #1

Answer

MC = ATC 2700 ATC=C_30° +2700 22016 1 =3q + 4600 3q+ = 69 3q² = 2700 q2 = 900 4= 30 P = ATC = 39 + 2700 =3* 30+ 2700 20 = 180

qE=30 units and P=180